From textbook:

"Let X have a negative binomial distribution with parameters r and p such that:

p(k) = p (1-p)*-, k=r,r+1,..

Find E[X] and Var[X] without using the definition; instead, consider how X can be written as a sum of independent random variables."

Question: How do I do this?
Answer #1

Here the pmf for ! Negative Binomial distribution is is f**)=P(K) = ) pr (1-P) KY; Kr, rt, . on Note that the negative BinomiTurm Tum Ely)= Ë ( 113.) p* (-2) ? Ž, (9*) ** (propos Typ Ž (721) pol (ap) ? 2-6 = (1-P) P XI or/1-P) P Note that the brack